OpenStudy (anonymous):

Add (-9y^2-9y)+(4y^2+2y-6)

OpenStudy (anonymous):

\[\left(4 y^2+2 y-6\right)+\left(-9 y^2-9 y\right)=-5 y^2-7 y-6 \]